In the daily production of a certain kind of rope, the number of defects per foot


Y is assumed to have a Poisson distribution with mean A = 2. The profit per foot


when the rope is sold is given by X, where X = 50 - 2Y - y2. Find the expected


profit per foot


1
Expert's answer
2022-05-23T14:07:26-0400

From the problem statement, "E[Y]=2" .

The Poisson distribution with parameter λ has mean λ and variance λ, so we have

"\lambda=2" and "V[Y]=2" . Now "E[Y^2]=V[Y]+E[Y]^2=2+2^2=6" .

Using linearity of expectation, 

"E[X]=E[50-2Y-Y^2]=E[50]-2E[Y]-E[Y^2]=50-2*2-6=40"